# STRIDE: Sparse Techniques for Regression in Deep Gaussian Processes

### v1

@@ 3.2 Maximisation step
$\frac{1}{S}\sum_{s=1}^{S}F_{V}\mathopen{}\mathclose{{}\left(\mathcal{M};\theta^{(s)}_{t+1}}\right)\eqqcolon F_{t}\mathopen{}\mathclose{{}\left(\mathcal{M}}\right)\>.$
Finding the global minimiser of $F_{t}$ likely requires an enumeration of $\{\mathcal{M}\subseteq[n]:|\mathcal{M}|=m\}$ which is infeasible. Instead, we follow a heuristic greedy strategy. We select the initial set of inducing points $\mathcal{M}_{0}$ in the algorithm using a strategy motivated by [38]: starting from an empty set, we add the observation that leads to the biggest increase of the objective. In practice, one chooses these observations from sets of candidate points $\mathcal{J}$ with cardinality $|\mathcal{J}|\ll n$ that are sampled randomly at each step to preserve computational efficiency. Elements are added in this greedy fashion until the maximum cardinality $m$ is reached.
When updating an existing set $\mathcal{M}_{t}$ to obtain $\mathcal{M}_{t+1}$ in STRIDE, we proceed as follows. We loop over every element in $\mathcal{M}_{t}$ and measure its contribution to the overall objective value by excluding it from $\mathcal{M}_{t}$ and evaluating $F_{t}$. In this way, we determine the element that, through exclusion, decreases $F_{t}$ the least, and exclude it from $\mathcal{M}_{t}$. We then loop over all observations in a set of candidates $\mathcal{J}\subset[n]\backslash\mathcal{M}_{t}$ to determine the element that, through adding, increases $F_{t}$ the most. We then include this element in our set of inducing points. This procedure can be repeated for a number of iterations. We choose the resulting set of inducing points as our updated set, $\mathcal{M}_{t+1}$. As mentioned, the set $\mathcal{J}$ is generally chosen to be of fixed cardinality $|\mathcal{J}|\ll n$ to control the cost of the update. As a result, the number of evaluations of $F_{t}$ needed for the update step is $R\,(m+|\mathcal{J}|)$, where $R$ is the number of update step repetitions.
@@ 3.3 Linear complexity on lower layers
The main advantage of using sparse GPR methods such as [38] is that they can be computed with an $\mathcal{O}(n)$ cost, as opposed to $\mathcal{O}(n^{3})$ for standard GPR. We aim at preserving this property in STRIDE. However, in order to be able to choose $\mathcal{J}$ as a subset of the observations, we need to evaluate the hidden layers of the deep GP at every observed location. This leads to having to construct covariance matrices of size $n\times n$, and sampling the hidden layers to obtain the covariance of the top layer then results in a computational cost of $\mathcal{O}(n^{3})$, which is the complexity of the required Cholesky decomposition. We now propose two approaches for efficient computation on the lower layers.
Adaptive cross approximation.
We can sample the hidden layers approximately using an approach that is cheaper than $\mathcal{O}(n^{3})$. In our experiments, we use an adaptive cross-approximation (ACA) algorithm [17, 18, 21], which approximates the Cholesky factor of the covariance matrix. More specifically, ACA determines a set of indices that identify the columns of the covariance matrix that can be used for an accurate low-rank approximation of the same covariance matrix. One can either select this set of indices once at the start of our EM algorithm and leave it fixed throughout, or, after each maximisation step (M), recompute the set of indices to reflect that the covariance matrix we aim to approximate has changed. In our experiments using ACA, we employ this latter approach. However, note that, since we are sampling whitened variables in the expectation step (E), this means that we also need to update these sampled variables to ensure they represent the same hidden parameters $f_{L-1:0}$ after updating our approximation of the covariance.
@@
We can sample the hidden layers at the current inducing points only. In this case, sampling the hidden layers becomes very cheap. The covariance at the top layer is then based on an interpolation of the next-lower layer to all observation locations. We perform this interpolation using standard GPR. The covariance kernel used here constitutes another set of tuning parameters. In practice, we simply use a stationary GP with the same marginal variance, correlation length and observation noise variance that the deep GP itself is initialised with. Using this “fully sparse” approach of sampling only at the inducing point locations also means that we need to update our hidden parameters every time we change the inducing points $\mathcal{M}_{t}$. That is, once the inducing points change, we now need evaluations of the hidden layers at these new locations. Once more, we obtain these through interpolation using GPR.
@@ 4 Numerical experiments
We now put our algorithm to the test in a number of real-life regression problems. In a first step, we consider the performance in three small-scale regression problems using datasets from the UCI Machine Learning Repository. [^1] These datasets can be found at https://archive.ics.uci.edu/. Since the deep GP can be specified in a multitude of ways, we use these examples to explore a range of possible setups. Mainly, we aim to investigate the effect of the deep GP construction type, as well as number of hidden layers. In a second experiment, we test the performance of our approach on a variation of the FashionMNIST dataset [42]. Namely, we consider features extracted using an auto-encoder, as presented in [5]. [^2] The dataset used is available at https://github.com/jwcalder/GraphLearning/blob/master/kNNData/FashionMNIST_vae.npz. This dataset is much larger (70,000 points with 30 dimensions) than the UCI datasets considered, and as such poses a significant computational challenge. Most importantly, standard GPR becomes infeasible on such large datasets, so that we have to resort to sparse GPR methods. All experiments were performed on an AMD EPYC 7513 32-core processor; we allocated 16 GB of RAM.
@@ 4.1 Standard UCI datasets
For our first experiment, we consider the energy efficiency, the concrete compressive strength, and the Boston housing datasets from the UCI Machine Learning Repository. These are relatively small, with numbers of data points ranging from 506 to 992. As a result, we can evaluate standard GPR on these datasets, and use the results for comparison. As we are using their strategy for selecting inducing points, we also compare with the sparse GPR approach from [38]. Furthermore, the datasets are small enough to run deep GPR without making use of inducing points. We refer to this as “standard” deep GPR. Both standard GPR and standard deep GPR would quickly become infeasible for larger datasets, since their computational cost grows cubically.
The goal of our experiment is to get an overview of the effect of choosing different architectures and approximation methods of the hidden layers. Our setup is as follows. The kernel $k$ is chosen as the Gaussian kernel, also known as the squared exponential or RBF kernel. The marginal variance, correlation length and observation noise variance are estimated through likelihood maximisation for the standard GP, and through maximising $F_{V}$ as given in eq. 2 for the sparse GP approach. The optimal hyper-parameters found in sparse GPR are then used to set up our deep GP model. That is, we use the same noise variance, and set the marginal variance of the top layer to the marginal variance obtained through optimisation. The marginal variances of the hidden layers are set to 1. The hidden layers $f^{(s,0)}_{L-1:0}$ are initialised so that the resulting covariance of the top layer is stationary with the correlation length found through hyper-parameter optimisation for each of the particles, $s=1,\ldots,S$.
@@
We use ten-fold cross-validation on each of the datasets to evaluate the errors of our individual GP and deep GP setups. Namely, we compute the standardised mean square error (SMSE; smaller values are better) as defined in [40]. The results are shown in Figure 2. In the left two plots, we obtain the SMSE for the energy efficiency dataset, using two and three layers in the deep GP models, respectively. In the third plot from the left, we show SMSE values for the concrete compressive strength dataset and in the plot on the right we consider the Boston housing dataset. For these last two plots, we use two layers in the deep GP models. We give the mean of the error values over the ten training/test data splits, as well as error bars of twice the standard deviation. The deep GP architectures are given in abbreviated form (“comp.” and “conv.”). The use of our two approaches to approximating hidden layers is marked by either “ACA” or “sparse”.
@@
Generally speaking, the standard deviations shown around the mean errors are especially large in the plot concerning the Boston housing dataset, as the regression results varied significantly between different splits of the data. However, we can still observe that the deep GP models appear to represent the data more accurately than the stationary GP models. When comparing the two different deep GP constructions, the composition architecture results in a better prediction error for each of the datasets. A reason could be that the convolution architecture makes use of GP layers with a one-dimensional output only, so that it is less flexible when storing information on the hidden layers compared to the composition architecture. We further note that using three instead of two layers improves the regression result for the energy efficiency dataset. The reduction in error can be substantial and can be observed for each of the deep GP setups.
@@
4.2 Large dataset (Fashion MNIST)
We now turn towards performing regression on a more challenging dataset. Fashion MNIST is a popular benchmark for machine learning methods, containing 70,000 greyscale images, each associated with a class label represented by an integer number [42]. As the images are of dimension 28$\times$28, the computational burden of using a deep GP with the composition architecture would make regression infeasible. However, the dimension of the data points can be reduced by techniques such as variational auto-encoders (VAEs) [5]. In our case, we use a variant of the dataset where the data points have already been transformed by such a VAE.
We want to explore the effect of the number of inducing points, $m$. To this end, we run sparse GPR [38], as well as STRIDE using the composition architecture and ACA with rank 50 to approximate the hidden layer. In STRIDE, we now use $S=10$ Markov chains as well as $R=3$ individual updates of the inducing points in each maximisation step (M) only. The size of the candidate set is chosen as $|\mathcal{J}|=1000$. Otherwise, the setup remains the same as in the previous experiment in Section 4.1. As the training data we use a random selection of 90% of the dataset, leaving the remaining 10% as the test set. We repeat this experiments for $m$ ranging from 50 to 125. The obtained SMSE values are presented in Table 1. Notably, the reported errors decrease as expected as we increase $m$. Even for small numbers of $m$, we obtain a more accurate regression result when using the deep GP model. However, even for sparse GPR, the errors do not stagnate as we increase $m$, indicating that this dataset requires large numbers of inducing points for an accurate representation.
Table 1: SMSE values obtained for different values of $m$.
$m$ 50 75 100 125
Sparse GP 0.2358 0.2360 0.2246 0.2169
STRIDE 0.2347 0.2219 0.2116 0.2063

### v2

@@ 3.2 Maximisation step
$\frac{1}{S}\sum_{s=1}^{S}F_{V}\mathopen{}\mathclose{{\left(\mathcal{M};f_{L-1:0}^{(s,t+1)}}}\right)\eqqcolon F_{t}\mathopen{}\mathclose{{\left(\mathcal{M}}}\right)\>.$
Finding the global maximiser of $F_{t}$ likely requires an enumeration of $\{\mathcal{M}\subseteq[n]:|\mathcal{M}|=m\}$ which is infeasible. Instead, we follow a heuristic greedy strategy. We select the initial set of inducing points $\mathcal{M}_{0}$ in the algorithm using a strategy motivated by [titsias2009variational]: starting from an empty set, we add the observation that leads to the biggest increase of the objective. In practice, one chooses these observations from sets of candidate points $\mathcal{J}$ with cardinality $|\mathcal{J}|\ll n$ that are sampled randomly at each step to preserve computational efficiency. Elements are added in this greedy fashion until the maximum cardinality $m$ is reached.
[…]
Input: Inducing points $\mathcal{M}_{t}$, observations $\mathcal{M}^{\ast}$, objective $F_{t}$, cardinality $J$, number of steps $R$.
Set $\overline{\mathcal{M}}\leftarrow\mathcal{M}_{t}$.
for $r=0,~1,~\ldots,~R-1$ do
 // Find point with least contribution.
   Set $\overline{F}\leftarrow$-$\infty$.
 for $z\in\overline{\mathcal{M}}$ do
[…]
         Set $\overline{z}\leftarrow z$.
         Set $\overline{F}\leftarrow F_{t}(\overline{\mathcal{M}}\backslash\{z\})$.
[…]
 // Replace with point with largest contribution.
   Set $F^{\ast}=$-$\infty$.
   Randomly draw candidate points $\mathcal{J}\subset\mathcal{M}^{\ast}\backslash\overline{\mathcal{M}}$ with $|\mathcal{J}|=J$.
 for $z\in\mathcal{J}$ do
    if $F_{t}(\overline{\mathcal{M}}\cup\{z\})>F^{\ast}$ then
         Set $z^{\ast}\leftarrow z$.
         Set $F^{\ast}\leftarrow F_{t}(\overline{\mathcal{M}}\cup\{z\})$.
  Set $\overline{\mathcal{M}}\leftarrow\overline{\mathcal{M}}\cup\{z^{\ast}\}$.
Set $\mathcal{M}_{t+1}\leftarrow\overline{\mathcal{M}}$.
Output: Updated inducing points $\mathcal{M}_{t+1}$.
Algorithm 2 Birth/death algorithm.
@@ 3.3 Linear complexity on lower layers
The main advantage of using sparse GPR methods such as [titsias2009variational] is that they can be computed with an $\mathcal{O}(n)$ cost, as opposed to $\mathcal{O}(n^{3})$ for standard GPR. More precisely, we consider the computational cost during training (offline phase), the computational cost of inference (online phase), as well as the memory consumption here. Asymptotically, sparse GPR as described in [titsias2009variational] has an online computational cost of $\mathcal{O}(nm^{2})$, an offline computational cost of $\mathcal{O}(nm^{3}|\mathcal{J}|)$, as well as a memory cost of $\mathcal{O}(nm)$.
We aim at achieving comparable properties in STRIDE. However, in order to be able to choose $\mathcal{J}$ as a subset of the observations, we need to evaluate the hidden layers of the deep GP at every observed location. This leads to having to construct covariance matrices of size $n\times n$, and sampling the hidden layers to obtain the covariance of the top layer then results in a computational cost of $\mathcal{O}(n^{3})$, which is the complexity of the required Cholesky decomposition. We now propose two approaches for efficient computation on the lower layers.
Adaptive cross-approximation.
We can sample the hidden layers approximately using an approach that is cheaper than $\mathcal{O}(n^{3})$. In our experiments, we use an adaptive cross-approximation (ACA) algorithm [harbrecht2012lowrank, harbrecht2015efficient, kressner20], which approximates the Cholesky factor of the covariance matrix. More specifically, ACA determines a set of indices $I$ that identify the columns of the covariance matrix that can be used for an accurate low-rank approximation of the same covariance matrix. More precisely, let $K\in\mathbb{R}^{n\times n}$ be a covariance matrix. Then, we denote by $K_{n,I}$ the matrix formed of the columns of $K$ at indices given by $I$. We furthermore denote by $K_{I,I}$ the matrix formed of the rows of $K_{n,I}$ at indices $I$. The low-rank approximation to $K$ induced by the ACA indices $I$ is then given by $K\approx K_{n,I}K_{I,I}^{-1}K_{n,I}^{T}$. We employ the ACA algorithm as presented in [kressner20] to compute a set of observation points, represented by indices $I$, that result in a low-rank approximation of the covariance matrix at the top layer. We use the same set of indices to approximate the covariance matrices at the hidden layers. Now, in order to obtain an approximate sample of the hidden layers, recall that our MCMC routine samples centred representations of the hidden layers, which we denote by $\xi_{L-1:0}^{(s,t)}$ here. The actual size of $\xi_{L-1:0}^{(s,t)}$ will depend on the rank $|I|$ of our approximation. We present the routine used for obtaining approximate samples $\hat{f}_{L-1:0}^{(s,t)}$ from the centred variable samples $\xi_{L-1:0}^{(s,t)}$ in Algorithm˜3.
One can either select this set of indices once at the start of our EM algorithm and leave it fixed throughout, or, after each maximisation step (M), recompute the set of indices to reflect that the covariance matrix we aim to approximate has changed. In our experiments using ACA, we employ this latter approach. However, note that, since we are sampling whitened variables in the expectation step (E), this means that we also need to update these sampled variables to ensure they represent the same hidden parameters $f_{L-1:0}$ after updating our approximation of the covariance.
Input: Centred MCMC sample $\xi_{L-1:0}^{(s,t)}$, pre-computed ACA indices $I$.
for $\ell=0,~\ldots,~L-1$ do
 if $\ell=0$ then
      Set covariance operator $k_{\ell}$ to stationary bottom layer kernel.
 else
      Evaluate covariance operator $k_{\ell}\mathopen{}\mathclose{{\left(\cdot,\cdot\,\Big|\,\hat{f}_{\ell-1}^{(s,t)}}}\right)$ w.r.t. lower layer approximation $\hat{f}_{\ell-1}^{(s,t)}$.
   Use $k_{\ell}$ to compute covariance matrices $K_{n,I}$ and $K_{I,I}$.
   Compute $U$ as the upper Cholesky factor of $K_{I,I}$.
   Evaluate approximate hidden layer sample $\hat{f}_{\ell}^{(s,t)}=K_{n,I}\,U^{-1}\,\xi_{\ell}^{(s,t)}$.
Output: Approximate hidden parameters sample $\hat{f}_{L-1:0}^{(s,t)}$.
Algorithm 3 Sampling with adaptive cross-approximation (ACA).
@@
We can sample the hidden layers at the current inducing points only. In this case, sampling the hidden layers becomes very cheap. The covariance at the top layer is then based on an interpolation of the next-lower layer to all observation locations. We perform this interpolation using standard GPR. The covariance kernel used here constitutes another set of tuning parameters. In practice, we simply use a stationary GP with the same marginal variance, correlation length and observation noise variance that the deep GP itself is initialised with. Using this fully sparse approach of sampling only at the inducing point locations also means that we need to update our hidden parameters every time we change the inducing points $\mathcal{M}_{t}$. That is, once the inducing points change, we now need evaluations of the hidden layers at these new locations. Once more, we obtain these through interpolation using GPR.
We summarise the fully sparse approach to sampling the hidden layers in Algorithm˜4. Here, we denote by $K_{m,m}$ the matrix of covariances between all points in $\mathcal{M}_{t}$, evaluated with the covariance kernel $k_{\ell}$. However, in order to be able to compute values of $F_{V}$, or evaluate the final deep GPR model given in Equation˜5, we require the last sampled layer, $\hat{f}_{L-1}^{(s,t)}$ to include function values not just at the inducing points $\mathcal{M}_{t}$, but at all observation locations. As a result, we include an extra interpolation step, in which we use stationary sparse GP regression to predict values of $\hat{f}_{L-1}^{(s,t)}$ at the missing locations.
Input: Centred MCMC sample $\xi_{L-1:0}^{(s,t)}$, set of inducing points $\mathcal{M}_{t}$.
for $\ell=0,~\ldots,~L-1$ do
 if $\ell=0$ then
      Set covariance operator $k_{\ell}$ to stationary bottom layer kernel.
 else
      Evaluate covariance operator $k_{\ell}\mathopen{}\mathclose{{\left(\cdot,\cdot\,\Big|\,\hat{f}_{\ell-1}^{(s,t)}}}\right)$ w.r.t. lower layer approximation $\hat{f}_{\ell-1}^{(s,t)}$.
   Use $k_{\ell}$ to compute covariance matrix $K_{m,m}$.
   Compute $C$ as the lower Cholesky factor of $K_{m,m}$.
   Evaluate approximate hidden layer sample $\hat{f}_{\ell}^{(s,t)}=C\,\xi_{\ell}^{(s,t)}$.
Interpolate $\hat{f}_{L-1}^{(s,t)}$ to all observation locations using sparse GPR.
Output: Approximate hidden parameters sample $\hat{f}_{L-1:0}^{(s,t)}$.
Algorithm 4 Fully sparse sampling approach.
3.4 Computational cost and memory consumption.
Using either of these two hidden layer approximations, we can now state the asymptotic costs associated with STRIDE. In the case of using an adaptive cross-approximation for the hidden layers, we assume that the ACA rank is chosen as a linear function of $m$. The computational cost of performing inference with the trained model is then given by $\mathcal{O}(nm^{2}S)$. That is, the online cost scales the same as for sparse GPR, with an additional dependence on the number of particles $S$. The computational cost of running STRIDE for training gives rise to a more complicated expression. Separated by expectation and maximisation steps, these costs are as follows,
combined cost of expectation steps (E): $\qquad\mathcal{O}(STK_{\mathrm{MCMC}}(nm^{2}+m^{3}(L-1)))\>,$
combined cost of maximisation steps (M): $\qquad\mathcal{O}(nm^{2}STR(m+|\mathcal{J}|))\>.$
Here, we denote the number of EM iterations by $T$, the number of MCMC steps used in the expectation step (E) by $K_{\mathrm{MCMC}}$, and the number of birth/death updates in the maximisation step (M) by $R$. Overall, we can observe that the computational cost remains linear in $n$.
The memory required for our approach still scales as $\mathcal{O}(nm)$ only. However, this is the minimum memory consumption and would increase for more parallel implementations. E.g., when fully parallelising the maximisation step (M), one requires memory scaling as $\mathcal{O}(nm|\mathcal{J}|)$.
@@ 4 Numerical experiments
We now test our algorithm in a number of numerical examples. In a first step, we explore the influence of the number of inducing points, $m$, and the number of layers, $L$, on the regression error. We perform this analysis in the setting of the one-dimensional toy problem already discussed in Section˜1. We then turn towards usual machine learning test problems. Namely, we consider the performance in three small-scale regression problems using datasets from the UCI Machine Learning Repository [boston_housing, concrete_compressive_strength_165, energy_efficiency_242]. These are benchmark datasets that allow us to compare our work to, e.g., [titsias2009variational, de2023thin, rossi2021sparse]. Since the deep GP in our approach can be specified in a multitude of ways, we explore a range of possible setups in these examples. Mainly, we aim to investigate the effect of the deep GP construction type, as well as number of hidden layers. In a final experiment, we test the performance of our approach on a variation of the Fashion-MNIST dataset [xiao2017fashion]. Here, we consider features extracted using an auto-encoder, as presented in [calder2020poisson, fashion_mnist_vae]. This dataset is much larger (70,000 points with 30 dimensions) than the UCI datasets considered. Importantly, standard GPR becomes infeasible on such large datasets, so that we have to resort to sparse GPR methods. All experiments were performed on an AMD EPYC 7513 32-core processor; we allocated 16 GB of RAM.
4.1 One-dimensional toy problem
As a first example, we consider the ground truth function already mentioned in Section˜1 and shown in the top left plot in Figure˜1. A similar function was also considered in [Osborne, roininen2019hyper]. In particular, the ground truth we consider here is defined on the interval $[0,1]$ and given by,
$f^{\dagger}(x)$ $=\begin{cases}\int_{0.1}^{0.4}\frac{1}{\sqrt{2\pi\>10^{-4}}}\mathrm{e}^{-\frac{(x-y)^{2}}{2\cdot 10^{-4}}}\sin\mathopen{}\mathclose{{\left(\frac{\pi(y-0.25)}{0.15}}}\right)\mathrm{d}y\>,&\text{if }0\leq x\leq 0.6\>,\\ -1\>,\hskip 0.0pt\hfill&\text{if }0.6<x\leq 0.75\>,\\ 1\>,\hskip 0.0pt\hfill&\text{if }0.75<x\leq 0.9\>,\\ 0\>,\hskip 0.0pt\hfill&\text{if }0.9<x\leq 1\>.\end{cases}$
Our goal is to use this as a simple example to compare STRIDE with standard GPR and the sparse GPR approach in [titsias2009variational], and to investigate the influence of the number of inducing points $m$ and the number of layers $L$ on the reconstruction result. We observe the ground truth function at $n=200$ locations chosen as a uniform grid in $[0,1]$, and pollute the observations with i.i.d. Gaussian noise with variance $\gamma^{2}=0.02^{2}$. As the ground truth contains jumps, we choose the kernel $k$ used for GP regression, sparse GPR, and to construct our deep GP prior, as the Matérn kernel with smoothness $\nu=3/2$ for its low regularity [williams2006gaussian]. That is,
$k(\mathbf{x},\mathbf{x}^{\prime})$ $=\sigma^{2}\mathopen{}\mathclose{{\left(1+\frac{\sqrt{3}\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}}{\lambda}}}\right)\exp\mathopen{}\mathclose{{\left(-\frac{\sqrt{3}\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}}{\lambda}}}\right)\>.$
Both for GPR and sparse GPR, we choose the marginal variance $\sigma^{2}$ and the correlation length $\lambda$ to maximise the likelihood of the observed data.
In order to define our deep GP prior to use in STRIDE, we employ a variation of the composition construction. Namely, we use the injective non-stationary kernel defined in eq.˜2, setting $u^{\mathrm{min}}=0.3$. In terms of parameters for STRIDE, we choose a total of 10 EM iterations, with each expectation step (E) consisting of 1000 MCMC steps, and each maximisation step (M) consisting of 5 inducing point update steps. The candidate set in the update steps, $\mathcal{J}$, is chosen as the entire observation data. As the number of particles $S$, we choose $S=10$.
First, we aim to explore the influence of the number of inducing points $m$ on the regression result. We select $m\in\{10,20,30,40,50\}$, and set the number of layers in the deep GP to $L=3$. We measure the quality of the regression results using the standardised mean square error (SMSE) as defined in [williams2006gaussian]. We repeat the experiment with 5 different random seeds, that is, with 5 different realisations of the observation noise. We present the results in the plot on the left in Figure˜3. We show lines for the mean SMSE of standard GPR (independent of $m$), sparse GPR, and STRIDE, with the shaded areas indicating a range of four ($\pm 2$) standard deviations around the mean error. We can observe that for sparse GPR, we need about 40 inducing points to reach the accuracy of GPR using the full observations. STRIDE, however, has the advantage of allowing for a non-stationary prior. As a result, we only require about 20 inducing points to reach an accuracy that is slightly better than standard GPR. Adding more inducing points then does not visibly affect the outcome of STRIDE.
In order to compare different numbers of layers in our deep GP model, we now set $m=20$. Otherwise, the setup of our experiment remains identical. We furthermore choose $L\in\{2,3,4,5\}$. The results of this experiment are presented in the plot on the right in Figure˜3. Again, we show lines for the mean SMSE of standard GPR (independent of $L$), sparse GPR (independent of $L$), and STRIDE, with the shaded areas representing a range of four standard deviations around the mean. We observe that the SMSE values obtained through STRIDE remain almost constant w.r.t. the number of layers $L$. That is, while adding one hidden layer seems to offer a slight advantage over standard GPR due to implementing a non-stationary prior, adding more layers does not seem improve the outcome in this example. This aligns well with the findings of [dunlop18] in their one-dimensional numerical experiment investigating the number of layers needed to accurately recover a non-smooth ground truth function.
Figure 3: Comparison of errors for different values of $m$ and numbers of layers $L$ in the 1D example.
@@ 4.2 Standard UCI datasets
For our first experiment, we now consider the energy efficiency, the concrete compressive strength, and the Boston housing datasets from the UCI Machine Learning Repository. These are relatively small, with numbers of data points ranging from 506 to 992. As a result, we can evaluate standard GPR on these datasets, and use the results for comparison. As we are using their strategy for selecting inducing points, we also compare with the sparse GPR approach from [titsias2009variational]. Furthermore, the datasets are small enough to run deep GPR without making use of inducing points. We refer to this as standard deep GPR. Both standard GPR and standard deep GPR would quickly become infeasible for larger datasets, since their computational cost grows cubically in $n$.
The goal of our experiment is to get an overview of the effect of choosing different architectures and approximation methods of the hidden layers. Our setup is as follows. The kernel $k$ is chosen as the Gaussian kernel, also known as the squared exponential or RBF kernel,
$k(\mathbf{x},\mathbf{x}^{\prime})$ $=\sigma^{2}\>\exp\mathopen{}\mathclose{{\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}}{2\lambda^{2}}}}\right)\>.$
The marginal variance $\sigma^{2}$, correlation length $\lambda$, and observation noise variance $\gamma^{2}$ are estimated through likelihood maximisation for the standard GP, and through maximising $F_{V}$ as given in eq.˜3 for the sparse GP approach. The optimal hyper-parameters found in sparse GPR are then used to set up our deep GP model. That is, we use the same noise variance, and set the marginal variance of the top layer to the marginal variance obtained through optimisation. The marginal variances of the hidden layers are set to 1. The hidden layers $f^{(s,0)}_{L-1:0}$ are initialised so that the resulting covariance of the top layer is stationary with the correlation length found through hyper-parameter optimisation for each of the particles, $s=1,\ldots,S$.
@@
We use ten-fold cross-validation on each of the datasets to evaluate the SMSE values of our individual GP and deep GP setups. The results are shown in Figure˜4. In the left two plots, we obtain the SMSE for the energy efficiency dataset, using two and three layers in the deep GP models, respectively. In the third plot from the left, we show SMSE values for the concrete compressive strength dataset and in the plot on the right we consider the Boston housing dataset. For these last two plots, we use two layers in the deep GP models. We give the mean of the error values over the ten training/test data splits, as well as error bars of twice the standard deviation. The deep GP architectures are given in abbreviated form (comp. and conv.). The use of our two approaches to approximating hidden layers is marked by either ACA or sparse. The setups that use STRIDE to perform regression are highlighted with a bold font.
@@
We use two-layer deep GPs in this experiment for the reason that the UCI datasets given here are known to be represented well by stationary Gaussian processes. As a result, we do not expect that adding more layers would improve the reconstruction results here. In fact, this is confirmed by our observation that the errors obtained for the energy efficiency dataset using two layers and three layers, respectively, appear very similar.
Generally speaking, the standard deviations shown around the mean errors are especially large in the plot concerning the Boston housing dataset, as the regression results varied significantly between different splits of the data. As a result, while the mean of the obtained errors makes it appear that the deep GP models represent the data more accurately than the stationary GP models, we cannot come to any conclusions here due to the variance of our results. When comparing the two different deep GP constructions, the composition architecture results in a better prediction error for each of the datasets. A reason could be that the convolution architecture makes use of GP layers with a one-dimensional output only, so that it is less flexible when storing information on the hidden layers compared to the composition architecture. We further note that using three instead of two layers improves the regression result for the energy efficiency dataset. The reduction in error can be substantial and can be observed for each of the deep GP setups.
As an alternative to the SMSE values presented here, we also report mean negative log-likelihood (MNLL) values for these experiments. To be precise, we compute the MNLL as,
$\mathrm{MNLL}=\frac{1}{2}\,\log\mathopen{}\mathclose{{\left(2\pi\gamma^{2}}}\right)+\frac{1}{2p}\,\mathopen{}\mathclose{{\left\|\mathbf{y}^{\mathrm{test}}-\mathbf{y}^{\mathrm{pred}}}}\right\|^{2}_{\gamma^{2}\mathrm{Id}_{p}}\>.$ (6)
Here, $\mathbf{y}^{\mathrm{test}}$ denotes the test data observations, and $\mathbf{y}^{\mathrm{pred}}$ is the prediction obtained through (potentially deep/sparse) GP regression. The integer $p$ represents the number of observations in our test data. Smaller MNLL values are generally better. As for the SMSE, we report mean MNLL values as well as a confidence intervals of two standard deviations computed by repeating the experiment for ten different training/test data splits. The corresponding error bar plots are presented in Figure˜5.
As is clear from the formula given in eq.˜6, the MNLL can be seen as a rescaling of the SMSE. As a result, our interpretation of the plots in Figure˜5 is almost identical with our discussion of the SMSE results. However, the choice of $\gamma^{2}$ is of importance here. In order to compute the MNLL, we use the same noise variance as used in our regression model. That is, $\gamma^{2}$ is computed via maximum likelihood estimation for standard GPR, and through maximising $F_{V}$ for sparse GPR. For STRIDE, we use the same noise variance as for sparse GPR. Since the estimated value of $\gamma^{2}$ was generally smaller for standard GPR compared to the other approaches, we observe a larger variance in the corresponding MNLL values than in the MNLL values obtained through sparse GPR and STRIDE.
Figure 5: Overview of obtained MNLL values in ten-fold cross-validation for three UCI datasets.
@@
4.3 Large dataset (Fashion-MNIST)
We now turn towards performing regression on a more challenging dataset. Fashion-MNIST is a popular benchmark for machine learning methods, containing 70,000 greyscale images, each associated with a class label represented by an integer number [xiao2017fashion]. As the images are of dimension $\times28$, the computational burden of using a deep GP with the composition architecture would make regression infeasible. However, the dimension of the data points can be reduced by techniques such as variational auto-encoders (VAEs) [calder2020poisson]. In our case, we use a variant of the dataset where the data points have already been transformed by such a VAE. Using such a transformation also has the advantage that it allows us to compare the data points, which represent images in this case, in the Euclidean distance.
We want to explore the effect of the number of inducing points, $m$. To this end, we run sparse GPR as described in [titsias2009variational], as well as STRIDE using the composition architecture and ACA with rank 50 to approximate the hidden layer. In STRIDE, we now use $S=10$ Markov chains as well as $R=3$ individual updates of the inducing points in each maximisation step (M) only. The size of the candidate set is chosen as $|\mathcal{J}|=1000$. Otherwise, the setup remains the same as in the previous experiment in Section˜4.2. As the training data we use a random selection of 90% of the dataset, leaving the remaining 10% as the test set. We repeat this experiments for $m$ ranging from 25 to 125. The obtained SMSE values are presented as an error bar plot in Figure˜7, with the exact numbers given in Figure˜7. For each value of $m$, we run the experiment five times, using a different train/test split of the dataset for each individual run. The values shown in Figure˜7 and Figure˜7 represent the average over five runs. The error bars in Figure˜7 cover $\pm 2$ standard deviations, while the values given in parentheses in Figure˜7 are the obtained standard deviations.
Notably, the reported errors decrease as expected as we increase $m$. Even for small numbers of $m$, we obtain a more accurate regression result when using the deep GP model. However, even for sparse GPR, the errors do not stagnate as we increase $m$, indicating that this dataset requires large numbers of inducing points for an accurate representation.
We present the computation timings obtained for this experiment in Figures˜11 and 11. They are split up into offline training times in the plot on the left and online inference times in the right plot. While sparse GP regression is significantly cheaper, computationally speaking, the cost of STRIDE scales similarly to sparse GP regression with increasing $m$. This is as we would expect from our discussion of the computational costs in Section˜3.4.
[…]